# The connection of polaritonic chemistry with the physics of a spin glass

### v1

@@ 2 Pauli-Fierz ab initio theory
To investigate the physical properties of the dressed electronic problem in Eq. (3), it is convenient to assume the dilute-gas limit for the corresponding many-electron wave function $\Psi$, i.e., free-space molecules do not interact with each other described by $H^{\rm m,e}(\boldsymbol{R})=\sum_{i=1}^{N}H^{\rm m,e}_{i}(\boldsymbol{R}_{i})$. This is a common choice in the field of polaritonic chemistry, which is applied to a broad range of different situations. 50, 51, 52, 18, 20 We note that this ansatz can also be extended to molecules in more complex environments, such as in solution. In this case $H_{i}^{\rm m,e}(\boldsymbol{R}_{i})$ does not correspond to the electronic structure of a single molecule but to a full solvation shell instead. Assuming non-overlapping electronic structures between the $N$ molecules, effectively reduces the total electronic wave function $\Psi$ to a simple Hartree product $\Psi=\psi_{1}\otimes\psi_{2}\otimes\dots\otimes\psi_{i}\otimes\dots\otimes\psi_{N}$ of $N$ single-molecule electronic wave functions $\psi_{i}$. These single-molecule wave functions are determined by the following $N$ coupled cavity Hartree (cH) equations, 20, 18
@@
Notice that the mean-field (direct product) description implies no quantum entanglement between the molecules. While in principle one could go beyond a mean-field theory, e.g., using coupled cluster, and other more accurate ab-initio methods53, 54, 55, 56, 57, 58, 37, 59, 60 for the collective electronic structure problem, it comes at the cost of increasing the computational load considerably. However, this additional complexity might not be necessary for large molecular ensembles under collective strong coupling, since it has been shown for similar situations that a mean-field treatment becomes exact in the large $N$-limit. 61, 62 However, the rigorous mathematical analysis of this aspect remains an open research question for the moment. Nevertheless, the subsequently developed connection to a classical spin glass is expected to sufficiently capture the most relevant physical mechanism of the inter-molecularly dressed electronic structure problem under VSC. Certainly, as always, the intra-molecular electronic structure problem (bare matter) requires the inclusion of exchange and correlation terms to be a chemically accurate description of a molecule. However, this is implicitly enabled by the cH equations. To determine the many-body ground state of the dressed electrons the $N$ coupled cH equations need to be solved iteratively until convergence. One immediately notices that the recursive dependency on $\sum_{j\neq i}^{N}\langle\psi_{j}|\hat{x}_{j}|\psi_{j}\rangle$ may introduce a significant (non-perturbative) modification of the electronic structure, even for small coupling constants $\lambda$61 This terms originates from the quadratic interaction term $\hat{x}^{2}$ in Eq. (1) that has the formal scaling of $N^{2}$. Notice, the length gauge representation is convenient to uncover this fundamental all-to-all interaction term. However, it is a gauge independent feature, which is implicitly present in any other gauge (e.g., also in the common velocity gauge). 57, 63 It makes the main difference to a free-space ensemble, where for each molecule the instantaneous electronic ground state can be determined independently and the unique electronic many-body ground-state is simple. In the free-space case the macroscopic ensemble has no influence on the individual molecule except of statistically exploring all single-molecule configurations. In the following, we will investigate the consequences of this novel long-range all-to-all interaction, which connects the properties of the macroscopic ensemble to its individual constituents.
3 A polarization glass is born
3.1 Collective strong coupling with the cavity Hartree equations
@@
To simulate an ensemble of molecules under VSC one has to propagate the classical equations of motion based on Eq. (5) in contact with a thermal bath. 5 Traditionally such a classical molecular dynamics setup would be considered overall in (canonical) thermal equilibrium. 64 However, thing are more intricate in a cavity 5, i.e., when forming a glassy phase as will be discussed in Sec. 4.2.4. For each propagation time-step the cH Eqs. (7) have to be solved self-consistently for all the electrons. This is numerically very expensive and demanding. Already minimizing the electronic structure problem of a single realistic molecule at a single time-step in principle poses a formidable challenge. 65 For this reason, a simple anharmonic one-dimensional Shin-Metiu model molecule 66 was used in Ref. 20 that can be solved numerically exactly and captures all relevant aspects to reach a qualitative understanding of an ensemble of molecules under collective VSC. The Shin-Metiu molecule is a paradigmatic and common model to study chemical reactions and conical intersections in- and outside of cavities 67, 68, 69, 12, 70. Each Shin-Metiu molecule consists of only one effective electron and nucleus and thus allows a numerically efficient solution of the free-space matter problem. This computationally simple model permitted the efficient exploration of classical cavity-Born-Oppenheimer molecular dynamics at finite temperature up to several 1000 molecules. 20 A first proof that strong collective coupling was reached in these simulations is displayed in Fig. 2, where the dashed blue line shows the collective Rabi splitting of the vibrational mode of the ensemble of Shin-metiu molecules. The two peaks correspond to the collective upper and lower polaritons, hybrid light-matter excitations, from which the field of polaritonic chemistry deduces its name. 3, 14, 7 The Rabi splitting is asymmetric, since the self-consistent solution captures the change in refractive index due to many molecules inside the optical cavity. 71, 72 For a purely harmonic electronic model one can show that this red-shift can directly be related to the static refractive index of the free-space electronic system. 72 Apart from these macroscopic properties, Fig. 2 also shows the averaged hypothetical absorption spectrum of a single molecule in the ensemble (solid blue line). Interestingly we find besides the dark states (vibrational states that decouple from the photon field) that also the local lower polariton is populated. In a next step, the cavity-induced polarization differences
$\Delta\mu_{0}=\langle\hat{\mu}_{i}\rangle_{0,\lambda=0}-\langle\hat{\mu}_{i}\rangle_{0,\lambda}$ (8)
@@
were analyzed in the collective electronic groundstate by increasing the number of molecules. However, at the same time the coupling was decreased as $\lambda/\sqrt{N}$ to see whether the local polarizations persist in the large-$N$ (thermodynamic) limit. The electronic polarization operator of the $i$-th Shin-Metiu molecule is given by $\hat{\mu}_{i}=-Z_{e}\hat{r}_{i}$. Indeed, Fig. 3 provides unique microscopic insights on the the scaling statistics of such cavity-induced polarizations. It depicts the first evidence of a cavity-induced local polarization under VSC in the collective-coupling regime. 20 In more detail, the time and ensemble averaged polarizations of the single molecules approach a finite value in the thermodynamic limit (blue). In contrast, the macroscopic polarization (black) quickly drops to a vanishingly small value. Notice, these findings hold qualitatively for an ensemble of aligned as well as randomly oriented molecules. 20 Naturally the question arises why do these small single-molecule polarizations not add up constructively to generate a macroscopic polarization? Here Fig. 3, which shows the time-averaged probability density $P_{\rm cH}(\Delta\mu_{0})$, details that overall the ensemble is not polarized, but there is a symmetric polarization ordering, i.e.,
$E_{\rm cH}[\Delta\mu_{0}]$ $=$ 00 0 (9)
${\rm Var}_{\rm cH}[\Delta\mu_{0}]$$\neq$ 0.0 0.0 . (10)
@@
This already vaguely resembles a spin-glass phase, for which, simply speaking, one observes zero overall magnetization, but ordering of the local spins (magnetization). 73, 74
Furthermore, we notice that the probability density appears relatively rough (almost discontinuous), even though it is an averaged quantity over many vibrational cycles (nuclear time scales). The emergence of such almost discontinuous probability distribution, as well as the very noisy data for the cavity-induced local polarizations (even on a logarithmic scale), suggest the presence of a frustration mechanism, i.e., the local polarization patterns are long-lived, despite being constantly perturbed by the vibrating nuclei. Moreover, the frustrated behaviour suggests that the time and ensemble averaged data is strongly correlated with the randomly chosen initial state of the system. This is what one would expect in a spin glass as well. 75 For this reason, the above local polarization pattern with zero net polarization was termed a polarization glass in Ref. 20. Eventually, we would like to highlight that solving the cH equations for the Shin-Metiu molecules numerically is only possible up to a certain collective Rabi-splitting. When increasing the collective coupling strength beyond this value, the self-consistency cycles of the cH equations do not converge anymore. 20 This suggests the emergence of a cavity-induced phase transition at a certain collective coupling strength. Again, the occurrence of a phase transition that depends on the strength of an all-to-all interaction term closely resembles the behaviour of a spin glass. In the following we want to further explore the connections of this frustrated polarizations in a cavity and the physics of spin glasses. We expect that this gives us some fundamental microscopic insights into the conundrum of VSC under collective strong coupling. But before, we discuss some experimental evidence for the proposed phase transition.

### v2

@@ 2.2 Dilute gas limit: cavity Hartree equations
Let us start with the simplest possible case to investigate the difference between free space and a cavity. For this we assume the dilute-gas limit for the corresponding many-electron wave function $\Psi$, i.e., free-space molecules do not interact with each other such that we can approximately replace $H^{\rm m,e}(\boldsymbol{R})\rightarrow\sum_{i=1}^{N}H^{\rm m,e}_{i}(\boldsymbol{R}_{i})$. Under this assumption one can determine the properties of a molecular ensemble from just solving a single representative molecule, which is the focus of usual quantum-chemistry methods 50. The gaseous ensemble properties are then determined by doing classical, independent statistics on top of the single-molecule spectrum. In this way one can connect to emission and absorption spectra of molecular ensembles. Indeed, these spectra are usually ensemble properties and the different peaks correspond to highly degenerate ensemble states. That is, there are very many combinations of single-molecule excitations that have the same total excitation energy of the ensemble. This simple setting is a common choice in the field of polaritonic chemistry, which is applied to a broad range of different situations. 57, 58, 59, 18, 20 Indeed, in principle we can extend this ansatz to molecules in more complex environments, such as in solution. In this case $H_{i}^{\rm m,e}(\boldsymbol{R}_{i})$ does not correspond to the electronic-structure of a single molecule but to a full solvation shell instead. And then classical, independent statistics are performed on top of this repeating unit. Remaining in the simple-dilute gas case, we essentially assume non-overlapping electronic-structures between the $N$ molecules, which reduces the total electronic wave function $\Psi$ to a simple Hartree product $\Psi=\psi_{1}\otimes\psi_{2}\otimes\dots\otimes\psi_{i}\otimes\dots\otimes\psi_{N}$ of $N$ single-molecule electronic wave functions $\psi_{i}$. The only difference to the free-space case comes in this level of approximation from the long-range $\hat{J}_{\rm DSE}$ term. Performing a minimzation with respect of the individual single-molecule wave functions (which internally still have all the intramolecular Coulomb and DSE contributions) are determined by the following $N$ coupled cavity Hartree (cH) equations, 20, 18
@@
Notice that the mean-field (direct product) description implies no quantum entanglement between the molecules. While in principle one could go beyond a mean-field theory, e.g., using coupled cluster, and other more accurate ab-initio methods60, 61, 62, 63, 64, 65, 37, 66, 67 for the collective electronic-structure problem, it comes at the cost of increasing the computational load considerably. However, as discussed above, if we are interested on the macroscopic scale the mean-field treatment (at least for the free-space case) becomes asymptotically exact, and we can expect reasonable results already on this level of mean-field theory. 54, 68 However, the rigorous mathematical analysis of this aspect when coupled to a multimode cavity remains an open research question for the moment. Nevertheless, the subsequently developed connection to a classical spin glass is expected to capture the most relevant physical mechanism of the dressed intermolecular electronic-structure problem under VSC. Certainly, the intramolecular electronic-structure problem (bare matter) requires the inclusion of exchange and correlation terms to be chemically accurate. However, since we do not specify the level of theory on the intramolecular scale in Eq. (10), this is implicitly enabled by the cavity Hartree equations.
To determine the dressed many-molecule ground state, the $N$ coupled cavity Hartree equations need to be solved iteratively until convergence. One immediately notices that the recursive dependency on $\sum_{j\neq i}^{N}\langle\psi_{j}|\hat{x}_{j}|\psi_{j}\rangle$ may introduce a significant (non-perturbative) modification of the electronic-structure, even for small coupling constants $\lambda$54 This terms originates from the quadratic interaction term $\hat{x}^{2}$ in Eq. (1) that has the formal scaling of $N^{2}$. Notice that the length-gauge representation is convenient to uncover this fundamental all-to-all interaction term. However, it is a gauge-independent feature, which is also present in any other gauge choice (e.g., also in the common velocity gauge). 64, 69 It is the main difference to a free-space ensemble, where the macroscopic state (which merely statistically explores all single-molecule configuartions) has no influence on the individual molecules. In the following, we will first investigate the consequences of this novel long-range all-to-all interaction on the individual constituents. Afterwards, we complement our picture by including cavity-mediated electron correlations in denser molecular ensembles under VSC.
2.3 Effective-electron approximation
To better understand the impact of the DSE Hartree interaction $\hat{J}_{\rm DSE}$ on the intermolecular electronic-structure of the molecular ensemble in a cavity, we additionally assume that every molecule has only one effective electron, i.e., setting $N_{e}=N$. This is a simple way of capturing the (electronic) polarizability of individual molecules, which allows to focus on the intermolecular properties of our ensemble. Those are assumed to not depend critically on the microscopic intramolecular details.
For one-dimensional and harmonic models of molecules, the self-consistent cavity Hartree problem in Eq. (10) can be solved analytically for arbitrarily many molecules $N$, as recently demonstrated in Ref. 70. In this simplified setting, we find an analytic expression for the feedback of the ensemble of molecules on the cavity mode, i.e., due to the change of refractive index 71, 70, in form of a renormalization of the cavity frequency $\tilde{\omega}_{\beta}$ (red-shift) as
$\tilde{\omega}_{\beta}^{2}$ $=$ $\gamma^{2}\omega_{\beta}^{2}$ (11)
$\gamma^{2}$ $=$ $\frac{1}{1+\lambda^{2}N\alpha_{i}}\leq 1.$ (12)
Here $\alpha_{i}$ corresponds to the bare matter polarizability of a single molecule and the redshift parameter $\gamma$ scales with the system size $N$ and thus depends on the collective coupling strength. The most notable analytic result is a cavity-modified local (!) molecular polarizability
$\tilde{\alpha}_{i}=\gamma^{2}\alpha_{i}<\alpha_{i}.$ (13)
Here the local molecular polarizability is defined by applying an external electric field $E_{\mathrm{ext}}$ to the full ensemble as 70
$\tilde{\alpha}_{i}=Z_{e}\frac{\partial\langle r_{i}\rangle}{\partial E_{\mathrm{ext}}}\ \mathrm{with}\ \hat{H}_{\rm tot}=\hat{H}^{e}-Z_{e}\sum_{i}^{N}\hat{r}_{i}E_{\mathrm{ext}}.$ (14)
Recent ab-initio calculation indicate that local polarizability modifications also exist for real molecular systems under collective strong coupling.72 In contrast, treating the external electric field with perturbation theory on the single-molecule level, where $\lambda$ seems a small parameter, the local polarizability of the molecule remains the free-space one, i.e.,
$\tilde{\alpha}_{i}^{\rm pert}\overset{N\gg 1}{\rightarrow}\alpha_{i}.$ (15)
It is easy to understand statistically why the usual trick of reducing to a single-molecule description, as is implicitly assumed in free space, does not work. The Hartree DSE term makes all the molecules statistically dependent, and hence we need to consider all possible cavity-mediated interactions between the molecules in the ensemble. From an energy perspective, we can equivalently say that while the prefactor $\lambda^{2}$ is small in $\langle\hat{J}_{\rm DSE}\rangle$, we have $N^{2}$ contributions that potentially have to be taken into account. This simple example shows that a self-consistent treatment of the full ensemble can become important to determine local properties under VSC.
2.3.1 Polarization glass
While the previous analytical results already show local effects under VSC for a externally driven system, we are interested specifically in the thermal equilibrium properties of molecular ensembles under VSC. To that end, one has to propagate the classical equations of motion from Eq. (5) in contact with a thermal bath. 5 To connect to the macroscopic experimental setup, one usually considers a fixed collective Rabi splitting (see, e.g., Fig. 2) and then slowly increases the number of molecules. That is, the macroscopic observable in absorption or transmission is the appearance of two peaks where there was one (highly degenerate) vibrational peak of the free-space ensemble. The distance between those two peaks is called the Rabi splitting and it is a measure of collective coupling strength $\lambda_{\rm coll}$ of the macroscopic ensemble. Having fixed the observed Rabi splitting, we perform a thermodynamic limiting procedure to see how different ensemble property behaves, and hence use $\lambda=\lambda_{\rm coll}/\sqrt{N}$ with $N\rightarrow\infty$. As indicated in Sec. 2, this limiting procedure should be taken with a grain of salt, since a cavity has a finite mode volume and the effective coupling strength is always non-zero. Thus, if the property does not converge for a given mode volume $\mathscr{V}$, then the presented level of description is most likely insufficient 38.
@@ 2.3.1 Polarization glass
Turning back to the harmonic model, the above procedure converges, however, to a rather boring result: The local effects vanish in the thermodynamic limit at thermal equilibrium 70. Of course, the setting is oversimplified and it is known that harmonic models often miss important physical aspects. So we consider a slightly more realistic example where the (effective) electron and nuclei do not interact harmonically. For this reason, an ensemble of an-harmonic Shin-Metiu model molecules 73 was used in Ref. 20. The Shin-Metiu molecule is a paradigmatic and common model to study chemical reactions and conical intersections in- and outside of cavities 74, 75, 76, 12, 77. This computationally simple model permits the efficient exploration of classical cavity-Born-Oppenheimer molecular dynamics at finite temperature up to several 1000 molecules. 20 In Fig. 2 we see the corresponding collective Rabi splitting (dashed blue line) for $N=900$. The lower peak is called the collective lower polariton and the upper peak the collective upper polariton, respectively. These peaks correspond to hybrid cavity-matter excitations, from which the field of polaritonic chemistry inherits its name. 3, 14, 7 The Rabi splitting is asymmetric, since the self-consistent solution leads to a change in refractive index (red-shift), as also discussed above for the harmonic model (see Eq. (11)). Apart from these macroscopic properties, Fig. 2 also shows the averaged hypothetical absorption spectrum of the average single molecule in the ensemble (solid blue line). Interestingly we find besides the dark states17, 20, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87 (vibrational states that decouple from the photon field) that also the local lower polariton is populated. As specific quantity that we want to consider in the above thermodynamic limiting procedure we consider the cavity-induced polarization differences
$\Delta\mu_{0}=\langle\hat{\mu}_{i}\rangle_{0,\lambda=0}-\langle\hat{\mu}_{i}\rangle_{0,\lambda}.$ (16)
@@
for increasing number of molecules. Here the electronic polarization operator of the $i$-th Shin-Metiu molecule is given by $\hat{\mu}_{i}=-Z_{e}\hat{r}_{i}$. Indeed, Fig. 3 is the first theoretical evidence of a cavity-induced local equilibrium polarization mechanism under collective VSC. 20 In more detail, the time and ensemble averaged polarizations of the single molecules approach a finite value in the thermodynamic limit (blue). In contrast, the macroscopic polarization (black) quickly drops to a vanishingly small value. Notice, these findings hold qualitatively for an ensemble of aligned as well as randomly oriented molecules. 20 The polarization pattern revealed by Fig. 3 can be summarized as
$E_{\rm cH}[\Delta\mu_{0}]$ $=$ $0,$ (17)
${\rm Var}_{\rm cH}[\Delta\mu_{0}]$$\neq$ $0,$ (18)
@@
which vaguely resembles a spin glass phase, for which, simply speaking, one observes zero overall magnetization, but ordering of the local spins (magnetization). 88, 89 For this reason, the above local polarization pattern with zero net polarization was termed a polarization glass in our previous work in Ref. 20.
Eventually, we would like to highlight that solving the cavity Hartree Eq. (10) for the Shin-Metiu molecules is only possible up to a certain collective Rabi splitting. When increasing the collective coupling strength beyond this value, the self-consistency cycles of the cavity Hartree equations do not converge anymore. 20 This suggests the emergence of a cavity-induced degeneracy/phase transition at a certain collective coupling strength. Again, the occurrence of a highly degenerate ground state that depends on the strength of an all-to-all interaction term closely resembles the behavior of a spin glass. In the following we want to further explore the resulting implications including electron correlation effects.